OpenStudy (anonymous):

limit of (sinn)^3/(3^n) as n goes to infinity

OpenStudy (zarkon):

\[\frac{-1}{3^n}\le\frac{\sin^3(n)}{3^n}\le \frac{1}{3^n}\]